### v2

# Quantifying the Dunkelflaute: An analysis of variable renewable energy droughts in Europe

Abstract
Variable renewable energy droughts, also called “Dunkelflaute”, emerge as a challenge for climate-neutral energy systems based on variable renewables. Drawing on 38 historic weather years and an advanced identification method, we characterize European drought events for on- and offshore wind power, solar photovoltaics, and renewable technology portfolios. We show that their characteristics heavily depend on the chosen drought threshold, questioning the usefulness of single-threshold analyses. Applying a multi-threshold framework, we quantify how the complementarity of wind and solar power temporally and spatially alleviates drought frequency, duration, and severity within (portfolio effect) and across countries (balancing effect). We identify the most extreme droughts and show how these drive major discharging periods of long-duration storage in a fully renewable European energy system, based on a policy-relevant decarbonization scenario. Such events comprise sequences of shorter droughts of varying severity. The most extreme event occurred in winter 1996/97 and lasted 55 days in a perfectly interconnected setting. While the average renewable availability during this period was still 47% of its long-run mean, we argue that system planners must consider such events when planning for storage and other flexibility technologies. Methodologically, we conclude that using single calendar years is not suitable for modeling weather-resilient energy scenarios.
CSP
Centralized Solar Power
ENTSO-e
European Network of Transmission System Operators for Electricity
DIETER
Dispatch and Investment Evaluation Tool with Endogenous Renewables
MBT
Mean Below Threshold
NTC
Net Transfer Capacity
PRL
Positive residual load
PV
photovoltaics
RoR
Run-of-River
TYNDP
Ten Year Network Development Plan
VRE
variable renewable energy
VMBT
variable-duration mean below threshold
keywords:
Variable renewable energy , Energy droughts
journal: arXiv
[…]
1 Introduction
Mitigating climate change requires deep reductions of greenhouse gas emissions that originate from fossil fuel use [lee_ipcc_2023]. Nuclear energy or fossil fuels in combination with carbon capture and storage are options for this [sepulveda_role_2018], but these firm low-carbon technologies face economic challenges and political controversies in many countries. Another major strategy for achieving climate neutrality is shifting energy supply to renewable energy sources, which are globally on the rise as they become increasingly cost-competitive [irena_renewable_2025]. However, the potential for dispatchable renewable energy sources such as hydro power, geothermal, or bioenergy is limited in most countries. Hence, future renewable energy systems are likely to heavily rely on variable wind and solar power [iea_net_2023, way_empirically_2022]. As these variable renewable energy (VRE) sources may become the most relevant primary energy sources in many countries [jacobson_100_2017, brown_response_2018, child_flexible_2019, pickering_diversity_2022, wang_accelerating_2023], energy systems are increasingly exposed to weather variability. In turn, spatial and temporal system flexibility is increasingly needed to match the supply of VRE sources with demand. This includes geographical balancing via transmission, different types of energy storage, and demand response [denholm_grid_2011, rasmussen_storage_2012, schlachtberger_benefits_2017, roth_geographical_2023, he_sector_2021, hunter_techno-economic_2021, tong_geophysical_2021].
While future energy systems with high VRE shares sources have to deal with variability on different timescales, extreme periods of low wind and solar availability emerge as a key challenge for realizing such renewable energy systems. Often referred to as “Dunkelflauten”, these VRE drought periods are characterized by long-lasting and substantial shortages of wind and solar energy [raynaud_energy_2018, kittel_measuring_2024] and may cover large geographical areas.If the availability of firm low-carbon generation technologies is low, dealing with VRE drought events necessitates the use of long-duration storage and other flexibility options [dowling_role_2020, sepulveda_design_2021]. Hence, understanding the spatial and temporal characteristics of VRE drought events is crucial for weather-resilient energy system planning and energy policy as well as for the design of energy markets and support instruments for generation and flexibility technologies. This includes questions as to how frequent, how long, and how severe such periods are, and assessing their spatial and temporal correlation across large-scale interconnected energy systems.
Literature from different fields contributes to VRE drought analysis. Wind droughts in the United Kingdom are well-studied [cannon_using_2015, potisomporn_spatial_2022, potisomporn_evaluating_2023, abdelaziz_assessing_2024], focusing on frequency-duration distributions, return periods as well as spatial and temporal correlations of historic or future on- and offshore wind droughts. Similar analyses for wind power have been conducted for Ireland [leahy_persistence_2013], the North Sea [patlakas_low_2017], Germany [ohlendorf_frequency_2020], or, analyzing deviations from climatological means, globally [antonini_identification_2024]. Additionally, research interest in drought patterns of policy-relevant portfolios comprising wind and solar photovoltaics (PV) is growing. Historic, future, and synthetic weather data of various world regions have been analyzed, such as Europe [raynaud_energy_2018, kies_critical_2021, kapica_potential_2024, hu_implications_2023, breyer_reflecting_2022, van_der_most_temporally_2024], the U.S. [rinaldi_wind_2021], China [li_renewable_2024], India [gangopadhyay_role_2022], Germany [kaspar_climatological_2019, mockert_meteorological_2023], Hungary [mayer_probabilistic_2023], Japan [ohba_climatology_2022], and Australia [richardson_climate_2023]. Besides regional and seasonal variations for single technologies and VRE portfolios, a general finding is that combining wind and solar within regions, as well as considering balancing across regions, can mitigate drought characteristics.
Different methodological approaches have been used for renewable drought identification [ohlendorf_frequency_2020, potisomporn_extreme_2024, kittel_measuring_2024, li_renewable_2024]. Droughts can be defined as periods of consecutive time steps with renewable availability below a certain drought threshold, either with fixed [raynaud_energy_2018, rinaldi_wind_2021, kies_critical_2021, gangopadhyay_role_2022, kapica_potential_2024, kaspar_climatological_2019, ohba_climatology_2022] or variable duration [potisomporn_extreme_2024], using various identification methods. For instance, these periods can be identified by searching for an availability constantly below the threshold [potisomporn_spatial_2022, potisomporn_evaluating_2023, abdelaziz_assessing_2024, ohlendorf_frequency_2020, kies_critical_2021, breyer_reflecting_2022, kaspar_climatological_2019, mayer_probabilistic_2023, ohba_climatology_2022], a mean availability over a certain averaging interval below the threshold [abdelaziz_assessing_2024, leahy_persistence_2013, patlakas_low_2017, ohlendorf_frequency_2020, raynaud_energy_2018, rinaldi_wind_2021, gangopadhyay_role_2022, kapica_potential_2024, mockert_meteorological_2023], or the deviation of the availability from a drought threshold [potisomporn_extreme_2024] or its climatological mean [antonini_identification_2024, stoop_climatological_2024]. Drought thresholds are either exogenously and presumably arbitrarily set [cannon_using_2015, potisomporn_extreme_2024, abdelaziz_assessing_2024, leahy_persistence_2013, patlakas_low_2017, ohlendorf_frequency_2020, kies_critical_2021, kaspar_climatological_2019, mockert_meteorological_2023, mayer_probabilistic_2023, ohba_climatology_2022], or derived from the data analyzed, such as a fraction of the time-invariant mean [raynaud_energy_2018, rinaldi_wind_2021, kapica_potential_2024] or maximum [breyer_reflecting_2022] availability, or of its time-variant climatological mean [gangopadhyay_role_2022, antonini_identification_2024, stoop_climatological_2024]. As each of these approaches has strengths and weaknesses, no standard drought identification method has yet emerged in the literature [kittel_measuring_2024].
Here we analyze and compare VRE drought events for single renewable technologies and for a fully renewable European electricity system, using VRE portfolio assumptions from a policy-relevant decarbonization scenario from European transmission grid operators. We do so for 33 individual European countries (EU27, the United Kingdom, Norway, Switzerland, and the Western Balkans), and for a pan-European “copperplate” scenario with perfect interconnection across all countries that allows for unconstrained geographical balancing. We draw on a large dataset of historical VRE availability factor time series covering 38 years [de_felice_entso-e_2022]. We use an advanced, open-source algorithm that fully captures unique drought periods, properly accounts for brief periods of higher renewable availability within longer drought events, and avoids arbitrary threshold choices [kittel_measuring_2024]. Our analysis illustrates the variety and seasonality of historical renewable energy drought patterns in Europe and provides evidence on how frequent, how long, and how severe such periods have been in the past, and to what extent they are correlated across the interconnected European energy system in space and time.
Our results quantify how the spatial and temporal complementarity of wind and solar PV alleviates VRE droughts within European countries, giving rise to a technology portfolio effect. Averaged over all thresholds and countries, the maximum drought duration of a renewable technology portfolio that combines solar PV as well as on- and offshore wind power decreases by 64%, 52%, or 47% compared to standalone PV, onshore wind, or offshore wind droughts. Our results also demonstrate how European integration can further mitigate VRE droughts by leveraging spatially complementary VRE availability profiles across regions, signified by a balancing effect. Considering unconstrained geographical balancing across all countries, the longest technology portfolio drought shortens by 65%. We further show that drought characteristics strongly depend on the chosen drought threshold and that single-threshold analyses lead to an incomplete characterization of extreme drought patterns. Using a multi-threshold analysis, we illustrate that the most extreme VRE events manifest as sequential droughts with varying severity, affecting multiple countries simultaneously to different extents. We introduce a drought mass indicator to identify the most extreme droughts and illustrate how these determine long-duration storage needs in a fully renewable European energy system. In Germany, the most extreme event occurred in the winter of 1995/96 and lasted 109 days. Under the assumption of unconstrained geographical balancing across Europe, the most extreme storage-defining event was substantially shorter and lasted 55 days, occurring in the winter of 1996/97. Such extreme droughts may occur at the turn of years, suggesting that planning horizons based on single calendar years are inappropriate for modeling weather-resilient future energy systems.
2 Results
We define a renewable energy drought as a period during which the moving average of the hourly VRE availability factor, also referred to as capacity factor, is below a given qualification threshold [kittel_measuring_2024]. As there is no consensus in the literature on appropriate drought thresholds, we apply a large number of thresholds to each investigated region and technology, ranging in 5% increments from 10% to 100% of the mean renewable availability across all investigated years [kittel_measuring_2024]. While low thresholds identify severe droughts, higher thresholds increasingly include more moderate drought events. By scaling drought thresholds relative to the mean availability factor of all hours in the data of each region-technology setting (henceforth referred to as “relative thresholds”), we account for regional and technology-specific differences in generation potentials. This enables meaningful cross-regional and cross-technological comparisons (for more details, see Section 4.1). To give an example, an identified wind power drought with a 24-hour duration at a threshold $\tau=0.1$ should be interpreted such that less than 10% of the long-run average electricity generation of wind power is available on average during a 24-hour period.
In the main part of the paper, we illustrate selected results for a limited set of countries. The Supplementary Information SI provides additional results. Interactive, high-resolution versions of all graphs that allow for retrieving concrete values for each of the illustrated 3D-plots as well as an additional animated graph are available online [kittel_high_2025].
2.1 Complex renewable energy drought dynamics and complementarities across regions and technologies
Figure 1 shows renewable energy droughts that last longer than one day for the years 1996 and 1997, which emerge to be particularly critical (see Section 2.5). The Figure combines these events for all employed relative thresholds, for selected countries from southern, central and northern Europe, as well as for the pan-European copperplate scenario. These countries represent some of the largest electricity markets in Europe, each with distinct yet complementary wind and solar generation profiles. The pan-European copperplate scenario combines respective profiles from all 33 countries in our data set, assuming unconstrained geographical balancing. The multi-threshold illustration shows the variety of renewable drought patterns in Europe across regions and technologies even within the same time period. It further indicates that the identified drought duration is highly sensitive to the underlying color-coded threshold. In general, single-threshold analyses using low thresholds can detect extreme short-duration droughts, which may occur in isolation or adjacent to high-availability periods (white spaces in the Figure). Higher thresholds increasingly identify longer-lasting events. These potentially include consecutive shorter periods of extreme renewable shortage or surplus, which are smoothed by the averaging mechanics of the identification algorithm (see Section 4.1). The multi-threshold perspective demonstrates that longer-lasting events identified by higher thresholds may encompass multiple severe droughts detected at lower thresholds.
Figure 1: Identified drought patterns in 1996 and 1997 for all investigated relative thresholds $\tau\in[0.1,1]$ and selected regions. For each technology-specific panel, a horizontal band indicates drought occurrences for the color-coded threshold of one region. To illustrate persistent patterns, only droughts lasting longer than one day are displayed. Figure SI.2 focuses on longer-lasting events by illustrating respective patterns for droughts lasting at least one week.
Solar PV events identified by moderate or high relative thresholds can last several months, indicating low solar availability during winter in Europe (Figure 1, panel b). Lower thresholds identify very severe PV droughts driven by low seasonal availability in combination with factors such as persistent cloud or snow coverage, such as in winter 1996. In summer, only rare and isolated PV droughts of lower severity occur that are longer than one day. These seasonal PV characteristics can generally be found across all years and are often more pronounced in Northern Europe, e.g., observable in winter 2011/12 (compare Germany or United Kingdom to Spain in Figure SI.3).
In contrast, severe long-lasting onshore and offshore wind droughts can occur throughout the year but tend to be more frequent in summer (Figure 1, panels c and d), which is in line with general wind seasonality [kaspar_climatological_2019]. The multi-threshold perspective further reveals that severe wind droughts identified by lower thresholds can be sequential and may occur within contiguous below-average wind periods detected at higher thresholds. The latter may last up to several months. While such long-lasting events usually occur in summer (compare Figures SI.3-SI.5), they may also take place in winter in some countries (e.g., in 1996/97, Figure 1). On- and offshore droughts are generally correlated, but may differ in severity (Figure 1, panels c and d).
The multi-threshold illustration also reveals a technology portfolio effect due to complementary wind and solar power droughts. Typically, high solar availability offsets wind droughts in summer, while wind power often compensates for the low solar PV availability in winter. This results not only in briefer but also less severe renewable portfolio droughts, i.e., higher thresholds apply in the same periods compared to single technologies (compare panel a of Figure 1 with panels b-d). However, if severe and persistent wind and PV droughts coincide, pronounced long-lasting compound droughts can occur even for renewable portfolios, as seen for instance in winter 1996/97.
The Figure also illustrates a geographical balancing effect. Assuming perfect interconnection across Europe, geographical balancing can substantially mitigate the severity and duration of renewable energy droughts. This can be seen when comparing the European copperplate scenario (top row in each panel) to the corresponding isolated-countries cases in Figure 1. This balancing effect is particularly pronounced for onshore wind, as onshore wind droughts tend to occur at different times across space, which allows for spatial smoothing across all 33 considered countries. The effect is less pronounced for offshore wind, which can only be deployed in countries with sea access. This limits the spatial extent over which balancing is possible. It is even less pronounced for solar PV because PV droughts are heavily driven by solar seasonality, which affects most countries simultaneously.
Geographical balancing also reduces the severity and duration of technology portfolio drought events as solar seasonality is lower in Southern Europe and wind droughts do not occur simultaneously across countries. Severe and long-lasting portfolio droughts may arise in single countries and also across Europe at the beginning of a calendar year, (e.g., observable in 2012, Figure SI.3), at the end of a year (e.g., 1982, Figure SI.4), or across the turn of years (e.g., in winter 1996/97, Figure 1). In some years, winter portfolio droughts are much less pronounced (e.g., 2013/14, Figure SI.5). Countries relying heavily on wind power, such as the United Kingdom, may also experience very severe events in summer.
While this section highlights important technology-specific and regional characteristics of renewable energy droughts in Europe drawing on specific weather years, it necessarily remains selective. In the following, we analyze the occurrence of European renewable droughts in a more systematic and exhaustive way for all technologies, regions, and weather years in our data, focusing on both shorter and very long-lasting events. In doing so, we also connect to and extend analyses based on single thresholds which have been prevalent in the literature so far.
2.2 Frequency-duration distribution: short drought events occur much more often than longer ones
A cumulative frequency-duration distribution shows how often VRE droughts that lasted at least a certain duration occurred on average per year in the investigated data. Figure 2 illustrates these distributions across all investigated relative thresholds for Germany and Spain, as examples of central and southern European countries with different wind and solar energy resources, and for the European copperplate scenario. For a given threshold, the frequencies of all events that are at least as long as a given duration are shown in ascending order on the vertical axis. For instance, at a relative threshold $\tau=0.75$, Germany experienced on average 9.8 onshore wind droughts lasting at least 48 hours (two days), 5.2 lasting at least 168 hours (one week), and 3.1 lasting at least 336 hours (two weeks) per year. Note that the duration axis in Figure 2 is truncated at 360 hours, so it focuses on the occurrence of shorter-duration events. These are particularly relevant for the operation of shorter-duration flexibility options in the power sector, e.g., battery or hydro reservoir storage. Longer-lasting drought exceeding 360 hours, which may drive the need for long-duration electricity storage or other firm low-carbon technologies, are not visible. The following sections focus more on such extreme periods.
Figure 2: Example of cumulative frequency-duration distributions of drought events across all investigated relative drought thresholds $\tau\in[0.1,1]$, sorting the yearly frequencies of all events that are at least as long as a given duration. White space indicates the absence of droughts for given thresholds in the data. The contour lines represent the threshold-specific yearly frequency. For illustration, the distributions are truncated at 360 hours, i.e., they show events with a maximum duration of just above two weeks. Figure SI.6 shows similar distributions for events lasting up to one full year. Yearly frequencies of events lasting at least two days, one week, and a fortnight are marked for a relative threshold of $\tau=0.75$.
In general, very short drought events occur much more often than longer ones for every given threshold. Events with very short duration reflect typical diurnal renewable variability, and should not be interpreted as exceptional periods of low availability. Long droughts lasting more than a week are infrequent. Importantly, the threshold choice substantially impacts the frequency of drought events. For lower thresholds, we generally find fewer events compared to higher thresholds. An exception is the left-hand side of the frequency-duration distributions shown in Figure 2. For very short drought durations, we find the largest number of events for relatively low thresholds, and a decreasing frequency of droughts for higher thresholds. This is because higher thresholds tend to identify fewer, yet longer-lasting droughts, which combine multiple shorter events that are counted as individual events for lower thresholds.
The cumulative frequency-duration distribution substantially differs between wind and solar power. For PV and low thresholds, we identify many events that last less than one day. They reflect the typical diurnal fluctuations of solar PV, with zero production at night-time [ueckerdt_analyzing_2015]. Increasing thresholds lead to fewer PV drought events, as many night-time periods of zero electricity generation merge into multi-day events. When the threshold approaches $\tau=1$, the identified droughts increasingly reflect a strong solar seasonality in Europe, with lower availability in winter than in summer. For on- and offshore wind power, droughts lasting less than one day occur less often. This is because wind power does not have regular non-availability at night-time as solar PV, but fluctuates less regularly. In turn, events that last longer than one day are more frequent, reflecting a typical multi-day variability of wind power [apt_spectrum_2007, ueckerdt_analyzing_2015].
The frequency-duration distribution of VRE portfolio droughts also illustrates the portfolio effect, i.e., droughts identified by lower thresholds (see white spaces in the portfolio row in Figure 2 are less frequent when wind and solar PV are combined due to complementary wind and solar availability. Hardly any VRE portfolio droughts arise for lower thresholds. For higher thresholds, their frequency still is much lower than for single technologies (compare highlighted frequencies in each column in Figure 2).
The Figure also illustrates the geographical balancing effect, as droughts occur less often in an ideal European interconnection than in single countries. This is true for both single VRE technologies and portfolios. The reason for this is that periods of low solar and wind availability are not perfectly correlated across Europe [roth_geographical_2023]. For example, assuming a relative threshold $\tau=0.75$, there have been on average approximately eight and five VRE portfolio droughts per year in Germany and Spain that lasted at least one week, respectively, while less than four of such events occurred in Europe.
In the SI, we provide additional results on the seasonality of drought patterns for Europe under the assumption of perfect interconnection. For low thresholds and events lasting longer than a few days, PV droughts are more frequent in winter, while wind droughts are more frequent in summer (Figure SI.7). The SI also shows return periods, which are the reciprocals of yearly frequencies. Similar to frequencies, return periods vary substantially between technologies and regions (see Section SI.4.)
2.3 Maximum drought duration strongly depends on threshold and differs across years and seasons
Figure 3 shows the longest droughts obtained from the data across all investigated years for different relative thresholds and technologies. A general finding is that the maximum drought duration strongly depends on the threshold. For very low thresholds around $\tau=0.1$, the maximum duration of all droughts in the data is very short. For higher thresholds approaching $\tau=1$, the maximum drought duration grows strongly and eventually reaches a plateau of 365 days, i.e., a full year. This is because at very high thresholds, the search algorithm identifies years with below-average renewable availability as drought events. Using very high thresholds close to $\tau=1$ thus appears not to be meaningful for identifying droughts relevant for energy system operations within a given year.
Figure 3: Most extreme duration of single drought events across all investigated years for given relative thresholds $\tau\in[0.1,1]$. The year with the most extreme event duration varies across thresholds. The marked numbers indicate the longest drought durations found for a relative threshold of $\tau=0.75$.
For medium thresholds, we find that both a technology portfolio and geographical balancing strongly decrease the maximum drought duration. For example, for a relative threshold of $\tau=0.75$, the longest PV drought in Germany lasts 291 days, while the maximum on- and offshore wind droughts are 303 and 209 days, respectively. In the case of a combined VRE portfolio, the respective maximum drought in Germany decreases to 106 days due to the technology portfolio effect. If, on top, perfect interconnection in Europe is considered, the duration of the longest portfolio event for the same threshold further declines to 55 days due to the geographical balancing effect.
Especially for medium thresholds, the maximum yearly drought duration strongly depends on the weather year (Figure SI.9). For example, the longest VRE portfolio drought in Germany obtained for a relative threshold of $\tau=0.75$ occurred in 1996 (106 days), while the shortest yearly maximum drought duration was in 2018 (22 days). Considering perfect interconnection across all countries and the same threshold, the longest and shortest portfolio droughts occurred in 1997 and 1999, respectively (55 and 8 days).
The most extreme droughts vary substantially across countries, both with respect to their duration and the year of occurrence (Figures SI.10 and SI.11). Importantly, the ranking of years varies with the drought threshold (Figure 3). For instance, the longest VRE portfolio drought in Germany for a threshold of $\tau=0.4$ occurred in 2007 but for the thresholds $\tau=0.6$ and $\tau=0.8$ the most extreme droughts were in 1997 and 2003. The corresponding longest drought events in Spain (or in Europe, assuming perfect interconnection) occurred in 1991 (2005), 2017 (1996), and 2017 (2002). The spreads between the longest drought events in each year further increase with higher thresholds, before they decrease again approaching the threshold $\tau=1$ (Figure SI.12). From these findings follows that any claim on a specific year being particularly relevant for renewable drought analyses should always be qualified with the respective threshold.
The maximum drought duration further varies substantially across seasons and technologies. Solar PV droughts primarily occur in fall and winter, with their median hour in November, December, or January (Figure 4). The maximum European PV drought under unconstrained geographical balancing lasts 248 days with a median hour in December for a threshold of $\tau=0.75$. The longest PV droughts with median hours in spring or summer months barely last longer than a few days, except for those identified by very high thresholds close or equal to $\tau=1$. In contrast, the longest European on- and offshore wind droughts mainly occur in spring and summer, with medians in July and maximum duration of 218 or 204 days, respectively. Maximum wind drought duration in winter months are much lower. In single countries, summer and winter wind droughts can be significantly longer (balancing effect). The seasonal complementarity of wind and solar further mitigates combined technology portfolio droughts to a substantial extent (portfolio effect). The longest-lasting droughts occur during winter months, which is a result of low seasonal PV availability in combination with rare but severe concurrent wind droughts. The median hours of the longest renewable technology portfolio droughts in Germany (106 days) or Spain (137 days) occur in November, or, under the assumption of unconstrained balancing, in January (55 days).
Figure 4: Example of most extreme duration of single drought events across all months in the data and all investigated thresholds $\tau\in[0.1,1]$. The contour lines represent the threshold-month-specific maximum duration. Note that the monthly assignment illustrates the median hour of respective droughts, while the duration of each event is plotted on the vertical axis. Events lasting longer than one month start (end) in previous (subsequent) months. The events with the highest and lowest duration across all months are marked for a threshold $\tau=0.75$.
2.4 Portfolio and balancing effects mitigate maximum drought durations
Comparing weighted averages over all thresholds and countries, the maximum drought duration of a VRE portfolio decreases by 64%, 52%, or 47% compared to solar PV, onshore wind, or offshore wind alone (Table 1). This portfolio effect is more pronounced for PV than for wind power, as onshore and offshore wind power complement both diurnal and seasonal shortages of PV. In contrast, wind power does not incur diurnal shortages. Assuming a European copperplate (CP), the portfolio effect (-80%) for solar PV is even stronger than for any isolated country. This is because in a larger geographical area, there is more complementary wind power available to compensate for solar PV shortages [roth_geographical_2023]. In turn, the portfolio effect is less pronounced for onshore wind power in the European interconnection (-45%) as PV availability is more homogeneously distributed. For offshore wind power, the portfolio effect is higher (-70%), as only a subset of countries features this technology, such that solar and wind onshore resources of all countries can compensate for extreme offshore droughts in a subset of countries.
Considering unconstrained geographical balancing across all countries, the longest PV, onshore wind, or offshore wind droughts decrease by 1%, 46%, or 34%, using weighted averages over all countries. The very small average balancing effect for PV, however, conceals that there is a substantial north-south divide between countries with higher and lower solar irradiation. For South European countries with abundant solar availability such as Spain, Portugal, Italy, or Greece, geographical balancing even increases solar PV drought duration. In turn, geographical balancing alleviates maximum PV drought durations in North European countries such as Germany or Scandinavia. In contrast, the balancing effect on wind drought duration is more similarly distributed across countries, as maximum wind droughts rarely occur simultaneously throughout Europe. The balancing effect is even more pronounced for VRE technology portfolios, decreasing by 65% on average. That is, geographical balancing not only helps to mitigate droughts for single renewable technologies, but even more so for a technology portfolio, due to complementary technology portfolios with different capacity mixes across countries.
Table 1: Change of the maximum drought duration as unweighted averages over all investigated thresholds $\tau\in[0.1,1.0]$ with $0.05$ increments in percent. The average portfolio and balancing effects are weighted averages considering the theoretical generation potential of each country.
Portfolio effect Balancing effect
VRE portfolio compared to single technologies European copperplate (CP) compared to single countries
region solar PV onshore wind offshore wind solar PV onshore wind offshore wind portfolio
AL -59 -58 19 -43 -64
AT -61 -46 -10 -38 -70
BA -11 -30 7 -52 -80
BE -71 -59 -48 -37 -57 -43 -70
BG -64 -69 34 -51 -58
CH -58 -55 -10 -51 -72
CZ -67 -51 -32 -50 -71
DE -77 -60 -54 -44 -53 -37 -65
DK -77 -60 -44 -45 -50 -26 -65
EE -68 -49 -31 -48 -58 -41 -77
ES -61 -58 -63 57 -34 -28 -53
FI -66 -32 -20 -50 -49 -34 -79
FR -72 -68 -66 -2 -48 -33 -50
GR -59 -38 -54 30 2 -27 -61
HR -51 -53 -5 -50 -72
HU -49 -28 -19 -39 -77
IE -53 -28 -5 -37 -48 -28 -79
IT -50 -59 -56 33 -52 -46 -68
LT -71 -43 -39 -46 -48 -37 -74
LU -69 -57 -39 -59 -72
LV -70 -53 -36 -47 -58 -37 -74
ME -58 -64 -7 -60 -72
MK -11 -11 20 -49 -81
MT 0 121 -78
NL -74 -59 -48 -42 -56 -41 -68
NO -65 -8 2 -46 -35 -18 -78
PL -66 -46 -37 -38 -52 -32 -73
PT -60 -56 -59 32 -31 -31 -62
RO -64 -58 1 -43 -62
RS -12 -23 6 -47 -79
SE -76 -53 -46 -47 -41 -24 -65
SI -22 2 -21 -55 -83
SK -49 -22 -27 -47 -79
UK -70 -59 -47 -37 -54 -35 -69
CP -80 -45 -70
Average -64 -52 -47 -1 -46 -34 -65
2.5 Drought mass: a multi-threshold metric to identify drought events relevant for long-duration electricity storage
Previous studies that use threshold-based identification methods typically use only a limited number of thresholds (e.g., [leahy_persistence_2013, cannon_using_2015, patlakas_low_2017, kaspar_climatological_2019, ohlendorf_frequency_2020, kies_critical_2021, potisomporn_extreme_2024, potisomporn_extreme_2024, abdelaziz_assessing_2024, mockert_meteorological_2023, mayer_probabilistic_2023, ohba_climatology_2022]). However, the findings presented above indicate that drought characteristics strongly depend on the chosen threshold. To account for this, we introduce a multi-threshold “drought mass” metric to identify the most extreme portfolio drought events and show that it identifies events that are relevant for long-duration electricity storage in renewable European energy systems. In contrast to previous single-threshold approaches, the drought mass quantifies extreme events by integrating both drought duration and severity across a wide range of thresholds, thus properly accounting for the complex and diverse renewable drought patterns discussed in Section 2.1. It accumulates the number of drought hours identified at thresholds $\tau\leq 0.75$ that occur within contiguous events identified at $\tau=0.75$, i.e., excluding droughts found by very high thresholds ($\tau>0.75$). The event with the highest score within a each pair of consecutive years in the data (i.e., 1982-1983, 1983-1984, 1984-1985, etc.) is identified as the most extreme drought event within this time frame. Section 4.2 further elaborates on the drought mass mechanics.
Figure 5 shows extreme portfolio drought patterns identified by the drought mass metric for the years 1996/97 and selected countries. The most extreme events, i.e., those with the highest drought mass, comprise sequences of shorter, but more severe droughts within contiguous well-below-average periods that may last up to several months. Events with the highest drought mass scores may occur in winter (teal boxes), potentially spanning across the turn of a calendar year, as in the case of the European copperplate in 1996/97. Severe drought mass events may also occur in summer (gray boxes) in countries that heavily rely on wind power such as the United Kingdom or Poland.
Note that renewable availability is not zero during these events, but it is low on average for a very prolonged period of time. For instance, the average availability factor of a renewable technology portfolio during the most extreme drought is 0.11 for the European copperplate scenario (55 days in winter of 1996/97), 0.07 in Germany (109 days in winter 1995/96), and 0.08 in Spain (131 days in winter 1988/89). For comparison, the average portfolio renewable availability factor over all hours in the data is 0.23 for the European copperplate, 0.19 for Germany, and 0.21 for Spain, respectively. That is, average renewable availability still amounts to 47% of its long-run average during the most extreme drought event that defines long-duration storage needs in a fully renewable, perfectly interconnected European power sector. For Germany or Spain, the average availability during the most extreme drought amounts to 37% or 35%, respectively. Yet, within the largest drought mass events, we find more severe but shorter drought events with very low VRE availability. For instance, during the extreme drought in the winter of 1996/97, events that lasted 17 and 18 days occurred in Germany with average availabilities of just over 0.05 identified by a threshold $\tau=0.5$, relating to 27% and 29% of the long-run average, respectively.
Figure 5: Most extreme drought events in 1996/97 of selected countries identified by the drought mass metric occurring in winter (teal boxes). For countries in which the most extreme drought events occur in summer, these are additionally shown (gray boxes). For each region, portfolio drought occurrences lasting longer than one day for color-coded thresholds (upper panel) as well as normalized exogenous demand profiles and optimized storage levels across three modeled interconnection scenarios (lower panel) are displayed.
Peak electricity demand often depends on ambient temperature and occurs in winter in most European countries, except for the Mediterranean area. Accordingly, winter drought events that coincide with high electricity demand are particularly relevant. Using a stylized electricity sector model (Section 4.3), we find that these compound events are a major driver for the use of long-duration electricity storage in all three interconnection states. Figure 5 shows the optimal operation of long-duration electricity storage for different degrees of interconnection between countries as well as electricity demand patterns for the years 1996/97. The most severe identified droughts largely coincide with the periods of long-duration storage discharge, i.e., decreasing storage levels. This holds true for both the European copperplate scenario and several isolated countries in the island scenario. When considering policy-relevant interconnection across all countries (“TYNDP”), the storage level patterns do not structurally change but only show a level shift compared to the island scenarios. That is, drought patterns identified for isolated countries are a reasonable approximation of drought patterns in settings with more policy-relevant, limited interconnection with regard to long-duration storage operation.
In some countries, the most severe drought mass event does not coincide with the major discharging period of long-duration storage. For example, in Sweden, Poland, or the United Kingdom, the metric finds that the largest drought events occur in summer (gray boxes in the lowest three panels of Figure 5). Long-duration storage usage, however, still coincides with the most extreme winter droughts, which are comparatively less severe. This is because electricity demand is lower during the more severe summer drought events identified in these countries. Consequently, summer droughts do not affect the system’s ability to meet demand as much. In complementary model runs where we assume a flat electricity demand profile, i.e., abstracting from seasonal and diurnal variability, the most severe event measured by the drought mass metric perfectly coincides with storage-defining periods in all countries, including Sweden, Poland, and the United Kingdom (Figure SI.13).
Figure 6: Drought mass of identified most extreme winter drought events. For each country or the European copperplate in the left panel, drought mass scores are normalized using the row-specific maximum. The colors of the right panel indicate the maximum duration of the event with the highest drought mass score normalized by the column-specific maximum, i.e., the maximum duration across all countries. The right panel also provides the year-pair with the most severe events as identified by the drought mass score per row and its corresponding duration in days.
The duration and severity of most extreme winter droughts captured by the drought mass metric varies significantly across years and countries (Figure 6). Assuming perfect interconnection between countries, the most extreme event in the data occurred in the winter of 1996/97 and lasted 55 days. This European super drought was caused by pronounced and partially overlapping droughts in several Central European countries and the United Kingdom (Figure 6). As the most pronounced events do not occur simultaneously in all countries, geographical balancing mitigates also such extreme droughts. Accordingly, the European super drought is substantially shorter than the most extreme droughts in nearly all isolated countries. Applying the drought mass metric to individual countries, we find the longest winter events in Eastern and Southern Europe. Further, smaller countries such as Slovenia (201 days, 1985/86) or Slovakia (195 days, 2015/16) tend to have longer extreme droughts than larger countries such as France (64 days, 2004/05), Sweden (83 days, 1997/98), Germany (109 days, 1995/96), or Spain (131 days, 1988/89). This is because smaller countries benefit less from geographical balancing within their borders.
2.6 Geographical balancing inside the European renewable energy super drought
Figure 7 shows four snapshots of selected hours within the European portfolio super drought, as identified by the drought mass metric at the turn of the years 1996/97. Each panel shows the drought severity present in the displayed hour for isolated countries (island scenario, left map) and for the European copperplate scenario (right map), illustrated by the lowest applying threshold. While the super drought affects multiple European countries simultaneously, the severity varies across countries and over time. Comparing the island scenario and the European copperplate shows that renewable energy shortages in the latter are always less severe than in the most affected countries at the same hour. This is because the assumed European interconnection mitigates shortages in one country by leveraging higher renewable availability in others. That is, geographical balancing is feasible even within the most severe pan-European drought event in the data. The graphs further illustrate that the super drought comprises sequences of shorter, but more severe droughts, both in individual countries (different colors on maps on the left-hand side) and on the European level (different colors between maps on the right-hand side). Further, even within the European super drought, there are hours with relatively high renewable availability (Figure 7d). The progression of sequentially occurring events inside the European super drought is even more visible in a complementary animated graph which is available as supplementary information [kittel_high_2025]. The animation, covering December 1996 and January 1997, shows that very severe droughts that cover most of Europe (e.g., visible in Figure 7a) are relatively brief and occur rarely, even inside the super drought event. That is, some renewable generation potential is always available somewhere in Europe. Even during the most extreme droughts, this can balance low renewable availability across technologies and regions to some extent.
(a) Severe droughts covering almost all of Europe on 7 December 1996 at 8pm.
(b) Severe droughts mainly in Central Europe on 19 December 1996 at 8am.
(c) Severe droughts covering almost all of Continental Europe on 10 January 1997 at midnight.
(d) Moderate droughts in most of Europe on 13 January 1997 at noon.
Figure 7: Snapshots of drought severity illustrated by the lowest applying threshold during the European super drought in winter 1996/97 for isolated countries (left) and under the assumption of unconstrained geographical balancing (right).
3 Discussion
Variable renewable energy droughts manifest in a wide range of patterns, from brief and isolated to very long-lasting, contiguous events of varying severity. Our multi-threshold analysis reveals that drought frequency, return periods, and duration are highly sensitive to the chosen threshold. We thus argue that a multi-threshold approach is required to adequately characterize heterogeneous and sequential patterns of renewable energy droughts. Single-threshold analyses, which are prevalent in the literature, are not capable of detecting such patterns and lead to an incomplete characterization of extreme droughts.
Based on a large data set and using a wide range of thresholds, we demonstrate that the complementarity of wind and solar power in Europe effectively reduces both the frequency and maximum duration of VRE droughts. Analyzing isolated countries, we find that combining solar and wind power leverages a portfolio effect, decreasing the maximum drought duration compared to single renewable technologies by on average 64% (PV), 52% (onshore wind), or 47% (offshore wind) in our European case study. This effect is particularly significant for PV, as wind power compensates for diurnal and seasonal shortages of solar energy. We further show that an unconstrained geographical balancing between European countries gives rise to a substantial balancing effect, which in our case study shortens the longest PV, onshore wind, offshore wind, or technology portfolio droughts by 1%, 46%, or 34%, or 65% on average. This effect is driven by an imperfect spatio-temporal correlation of extreme drought periods, especially concerning wind droughts.
To overcome the challenges of single-threshold analyses, we introduce a drought mass metric that enables an integrated assessment of the duration and severity of events. It reveals that extreme drought events of varying severity may emerge sequentially within contiguous low-availability periods lasting several weeks or months. These events affect multiple European countries simultaneously, yet with varying intensity. In most countries and in a perfectly interconnected Europe, the largest drought events occur in winter. In wind-dominated systems, however, the most severe events may occur during summer. We show that compound winter events, characterized by extreme technology portfolio droughts in combination with peak electricity demand, determine the major discharging period of long-duration electricity storage in a fully renewable European energy system.
The severity of such events varies substantially across years and countries, enabling geographical balancing to mitigate the need for system flexibility. Using the drought mass metric and drawing on renewable capacity assumptions from a policy-relevant scenario, we find that the most severe renewable technology portfolio drought observable in the data occurred in the winter of 1996/97, assuming a perfectly interconnected, fully renewable European energy system. This European super drought event lasted 55 days. The maximum droughts in individual, isolated countries are much longer, e.g., 109 days in Germany. This shows that even during the most pronounced European drought event, geographical balancing can be leveraged.
Previous research has highlighted the importance of diversifying variable renewable energy supply [heide_seasonal_2010, jurasz_review_2020, harrison-atlas_temporal_2022] and advancing the integration of European power systems via transmission grid expansion [schaber_parametric_2012, fursch_role_2013, roth_geographical_2023] for realizing cost-efficient renewable energy systems. Based on our analysis, we not only underscore this notion but extend it by arguing that technology portfolios and geographical balancing are also important strategies for effectively dealing with rare and extreme renewable energy droughts. We also show that the most severe VRE portfolio droughts largely drive the need for long-duration storage in energy systems fully based on variable renewable energy sources.
Our multi-threshold analysis facilitates a comprehensive characterization of renewable drought events in Europe, solely based on VRE availability time series. Our approach, and particularly the drought mass metric, is not prone to artifacts that may arise in drought analyses relying on energy system modeling such as the duration of the planning horizon or technology cost assumptions. These may substantially affect optimal long-duration storage investments [dowling_role_2020]. Yet, methodological challenges remain. Drought analyses purely based on renewable availability time series cannot take into account real-world interconnection levels, which are between our extreme assumptions in the island and copperplate scenarios. Drought characteristics of scenarios with cross-regional exchange are thus likely between these corner solutions. For instance, the duration of the European super drought is likely to last longer than the 55 days identified for the copperplate scenario, but shorter than identified here for isolated countries (Figure 6), depending on future interconnection capacities between countries.
Further, the drought mass indicator does not consider a range of aspects that energy system models do. For instance, it does not capture periods with very high renewable availability that may arise between or after extreme droughts. Energy models with perfect foresight optimize long-duration storage operations considering not only droughts but also high-availability periods. This may result in major storage discharge periods that occur at different times or last longer than the most extreme drought periods identified by the drought mass indicator. In addition, the drought mass does not account for storage conversion losses or renewable curtailment. The latter may be required even within longer-lasting droughts in case of brief periods with very high availability that exceed storage charging capacities [ruhnau_storage_2022]. It is possible to integrate conversion losses into drought identification methods, but this comes at the cost of higher complexity [kittel_measuring_2024].
Due to computational limits, energy system models are often solved for a limited number of weather years, or even for only a single weather year. While substantial inter-annual variation in optimal energy model outcomes has been documented in the literature [pfenninger_long-term_2016, staffell_using_2016, pfenninger_dealing_2017, zeyringer_designing_2018, collins_impacts_2018, schlott_impact_2018, kaspar_climatological_2019, hill_effects_2021, tong_geophysical_2021, grochowicz_intersecting_2023, gunn_spatial_2023, ruggles_planning_2024], our analysis shows that extreme portfolio droughts also vary significantly across different years and regions. This likely aggravates inter-annual variations of energy model outcomes, particularly in terms of long-duration system flexibility needs. Accordingly, it appears desirable to corroborate the findings and conclusions of previous model-based energy system studies that are based on a small number of weather years or even only a single one and do not include extreme VRE droughts. For example, the European Ten Year Network Development Plan 2022 draws on the single weather year 2009 [entso-e_tyndp_2022]. Long-term energy scenarios developed for the European Commission generally lack transparency about actual weather years used, but appear to be based on a very limited number of weather years [european_commission_directorate_general_for_energy_eu_2021]. The same holds true for influential climate neutrality scenarios for Germany [bdi_klimapfade_2021, fraunhofer_isi_langfristszenarien_2024, prognos_klimaneutrales_2021, luderer_deutschland_2021]. Our analysis shows that the winter of 1996/97 including the pan-European super drought is particularly relevant for weather-resilient planning of the European energy system. It further indicates relevant weather years for single-country analysis.
Energy model planning horizons that are in line with a single calendar year further appear unsuitable for investigating weather-resilient future scenarios as extreme, storage-defining drought events may extend across the turn of years. Instead, we recommend adopting a planning horizon that captures relevant renewable seasonality patterns. For European settings, this may require at least bi-annual or single-year summer-to-summer planning horizons. Multi-annual planning horizons would be preferable but may come with increasing numerical challenges [grochowicz_using_2024, ruggles_planning_2024], and the length of the model period can impact optimal storage capacities [dowling_role_2020].
Additionally, multi-sectoral and technology-rich energy system models with detailed grid representations often have to use aggregated temporal resolutions to reduce the computational burden [nahmmacher_carpe_2016, kittel_temporal_2022, teichgraeber_time-series_2022, goke_adequacy_2022]. Our results indicate that the dynamics of storage-defining extreme drought events should adequately be reflected by representative time slices used in such models [sanchez-perez_effect_2022]. This is even more relevant for integrated assessment models, which are also used for long-term energy system analyses [gong_bidirectional_2023, schreyer_distinct_2024], but usually come with even coarser time resolutions than energy system models.
Renewable drought analysis based on renewable availability time series, as conducted in this paper, is not intended for precisely quantifying long-duration storage requirements. Instead, it complements energy system modeling, which is required for evaluating flexibility needs in renewable energy systems for the reasons discussed above. Insights on extreme droughts can support the selection of relevant weather years for computationally intensive multi-carrier and multi-sectoral energy system models. Additionally, such insights can help explaining the outcomes of energy models, particularly concerning long-duration flexibility needs.
We see several promising avenues for future research. We note that a wide range of different datasets of renewable availability are used for energy system modeling, with sometimes conflicting model outcomes and policy recommendations [kies_critical_2021]. Assessing the sensitivity of VRE drought characteristics to such datasets could provide valuable feedback to the meteorological research community that generates and continuously improves such datasets [craig_overcoming_2022]. Further, the temporal extent of most data sets is limited to a few decades. Statistically robust findings for low-probability VRE drought events with high impact are likely to require longer records of weather data, which also might involve a need for synthetic datasets, e.g., applied in Gangopadhyay et al. [gangopadhyay_role_2022]. Moreover, drought analyses for modeled future weather data that incorporate the effects of global climate change would be desirable. Finally, characterizing renewable energy droughts in other world regions and comparing these with our findings for Europe would be of interest. While parameters like the cut-off threshold of the drought mass metric ought to be reviewed, we argue that our multi-threshold framework is particularly useful for respective analyses in world regions where renewable availability is characterized by diurnal or seasonal variability, e.g., related to regional weather phenomena such as monsoon events.
4 Methods
4.1 Variable renewable energy drought definition and identification
Our VRE drought analysis is based on hourly VRE availability factor time series. An hourly availability factor, also referred to as capacity factor, ranges between 0 and 1 and indicates the generation potential (in MWh) of a VRE source in a specific region and hour normalized by its capacity (in MW). We use a new open-source method that applies a variable-duration mean below threshold (VMBT) algorithm for drought identification. This method searches for periods with a moving average below a given drought qualification threshold by iteratively decreasing the event duration. In each iteration, the algorithm sets the averaging interval to the respective event duration, starting with very long-lasting events and iteratively continuing to such that last only a few hours. A moving average below the drought threshold identifies a drought event. It is excluded from subsequent iterations, in which the averaging interval decreases further and additional (shorter) events are identified. The algorithm is introduced and discussed in Kittel & Schill [kittel_measuring_2024].
This iterative procedure overcomes shortcomings of previous research [kittel_measuring_2024]. It pools adjacent periods that independently may not qualify as VRE drought, identifies unique events with the longest duration possible, avoids double counting as well as overlaps with adjacent events, and ensures that the full temporal extent of drought periods is captured [kittel_measuring_2024]. To cover the full spectrum of VRE droughts, we parameterize the VMBT algorithm to iteratively search for periods ranging in duration from two years to one hour in descending order. The code and input data of the drought analysis tool are freely available on GitLab at https://gitlab.com/diw-evu/variable_renewable_energy_droughts_analyzer.
While threshold-based renewable drought identification methods are commonly applied in different bodies of the academic literature [kittel_measuring_2024], previous threshold-based analyses of renewable droughts across multiple regions, technologies, or technology portfolios often apply uniform, exogenously defined thresholds e.g., [leahy_persistence_2013, cannon_using_2015, patlakas_low_2017, kaspar_climatological_2019, ohlendorf_frequency_2020, kies_critical_2021, potisomporn_extreme_2024, potisomporn_extreme_2024, abdelaziz_assessing_2024, mockert_meteorological_2023, mayer_probabilistic_2023, ohba_climatology_2022]). Such approaches implicitly assume that one unit of generation capacity has a uniform generation potential across all region-technology settings. However, substantial differences in annual generation potentials may prevail due to varying meteorological conditions, leading to comparability issues in the resulting drought characteristics. To address this, threshold-based drought analyses should account for cross-regional and cross-technological variations in renewable generation potential when defining drought thresholds. Thresholds scaled relative to the mean availability of each setting (termed here as “relative thresholds”), reflect these variations [kittel_measuring_2024].
Additionally, as highlighted in the literature review, no consensus has yet been established regarding the selection of concrete threshold values. On the contrary, they often appear arbitrarily chosen or lack clear justification [kittel_measuring_2024].
To address both the challenge of comparability across regions and technologies and appropriate threshold value selection, we apply a wide range of relative drought thresholds, each scaled relative to the multi-year mean availability of the respective time series, i.e., each single technologies or different VRE portfolios across countries. Threshold values $\tau$ range from 10% to 100% of the mean availability factor over all investigated years, increasing in 5% increments. This broad range allows us to capture a variety of renewable drought events. Very low thresholds capture potentially brief but very severe drought events with near-zero renewable availability (e.g., $\tau=0.1$), whereas high thresholds (e.g., $\tau=1.0$) are more likely to identify very long-lasting events or even full weather years with below-average renewable availability. Figure SI.1 illustrates the relationship between these relative thresholds and absolute mean availability factors.
We use country-level VRE availability time series from the Pan-European Climate Database provided by the European Network of Transmission System Operators for Electricity [de_felice_entso-e_2022], comprising 38 weather years from 1982 to 2019. These data have been derived from reanalysis data, which has been converted to renewable availability by the Transmission System Operators. The database has been used for policy-relevant strategic reports, such as the Ten Year Network Development Plan (TYNDP) 2022 or the European Resource Adequacy Assessment 2021. We use exogenous renewable capacity assumptions to generate the capacity-weighted VRE portfolio time series as explained in the next paragraph. These assumptions stem from the TYNDP 2022 (scenario Distributed Energy), which optimized these portfolios for a climate-neutral European power sector in 2050. For Germany, we update the capacity targets according to the latest government targets for 2045 [bundesamt_fur_justiz_gesetz_2024], in which Germany aims to achieve climate-neutrality.
Multi-regional VRE drought analyses based on VRE availability time series can be conducted based on two extreme assumptions of electricity transmission between countries [kittel_measuring_2024]: either perfect interconnection across all countries (“copper plate scenario”) or complete isolation of all countries (“island scenario”). For the former, all countries are treated as one single pan-European node. All regional time series are combined into a composite using capacity-weighted averages, with weights according to the capacity assumptions from TYNDP 2022. For the latter scenario, thresholds are scaled as discussed above. Accordingly, regional VRE portfolios are constructed using capacity-weighted averages of all contributing technologies.
4.2 Drought mass: a multi-threshold metric to identify storage-defining drought events
Here, we introduce a drought mass metric to identify storage-defining drought events in a renewable energy systems. This metric identifies and ranks extreme events by integrating both drought duration and severity across a wide range of thresholds. As discussed in Section 2.3, very high relative thresholds close to $\tau=1.0$ are not meaningful, as they tend to capture very extended periods or even whole weather years with below-average renewable availability, instead of unusual drought events. For the drought mass metric computation, we therefore limit our analysis to drought events identified at thresholds $\tau\in[0.1,0.15,...,0.75]$, a set with cardinality $c=14$. For each threshold $\tau$, we first generate binary time series for each year in the data, which assign the value $1$ to hours that qualify as drought and 00 otherwise. These time series are then paired into consecutive two-years periods (e.g., 1982-1983, 1983-1984, 1984-1985, etc.), resulting in vectors with a magnitude of $T=17520$ hourly time steps per pair. This results in 37 year-pairs and accordingly allows analyzing 37 complete winter droughts. Note that the investigated data abstract from leap days. Next, we stack binary threshold-specific vectors to form a matrix of dimensions $T\times c$, where each row corresponds to a specific hour and each column to a threshold. Within each two-year period, we search for contiguous drought events identified at $\tau=0.75$, starting at hour $k$ and ending at hour $l$. For each of these intervals $[k,l]$, we compute the total number of drought hours across considered thresholds, i.e., we sum across all $c$ columns. The event with the highest cumulative multi-threshold score is then selected as the most extreme drought event for each weather year pair.
The drought mass metric equally weighs drought events identified across all considered thresholds. We chose a cut-off threshold $\tau=0.75$, which defines the event duration. Within an contiguous event, all thresholds $\tau\leq 0.75$ equally contribute to an event’s overall score. We tested a wide range of alternative cut-offs and weighting schemes to identify the most effective drought mass design. The best alignment between events with the highest winter drought mass scores and major storage discharge events was achieved at a cut-off threshold of $\tau=0.75$ and a simple approach of equal weighting across all considered thresholds.
Countries with high shares of wind power in their capacity mix may experience the most extreme portfolio droughts in summer. Peak electricity demand periods often occur in winter, except in some South-European countries. To account for this, we compute a drought mass score for droughts occurring throughout the year and another one relating only to winter droughts excluding the period from May until September. When illustrating the relation of drought patterns and long-duration electricity storage use, we display both the most extreme summer and winter droughts if the highest drought mass score relates to summer droughts (compare regions with gray and teal boxes in Figure 5). Conversely, if the highest drought mass score throughout the year relates to a winter drought, we mark only one event (compare regions with a gray box only in Figure 5).
4.3 Power sector modeling and scenarios
We use the open-source power sector model Dispatch and Investment Evaluation Tool with Endogenous Renewables (DIETER) to analyze the interaction between VRE droughts and long-duration storage needs in a fully renewable European power sector. DIETER is a linear optimization model that determines least-cost capacity and dispatch decisions based on an hourly resolution [zerrahn2017, gaete-morales_dieterpy_2021]. Different versions of the model have been used to study various aspects of VRE integration and their interaction with other flexibility options or sector coupling technologies [zerrahn2018, schill_long-run_2018, schill_electricity_2020, schill_flexible_2020, kittel_renewable_2022, roth_geographical_2023, kirchem_power_2023, gong_bidirectional_2023, gaete-morales_power_2024, roth_power_2024]. Here, we use a model version that includes 33 European countries (EU27, the United Kingdom, Norway, Switzerland, and the Western Balkans). As very long-lasting VRE droughts may span across the turn of a calendar year, we extend DIETER’s planning horizon to two full years. For transparency and reproducibility, we provide the model code, the input data, and a manual in a public repository under permissive licenses available on GitLab at https://gitlab.com/diw-evu/projects/quantifying_the_dunkelflaute_energy_system_analysis.
We use policy-relevant renewable capacity mixes from the TYNDP 2022 (scenario “Distributed Energy”). Electricity demand profiles are derived from the European Resource Adequacy Assessment 2021 (target year 2025), including limited electrification of heat and transport. These profiles are scaled to the TYNDP demand levels in 2050, and adjusted for net-importing and net-exporting countries to ensure they can be met by domestic renewable supply. Inter-annual variations are mainly driven by temperature differences.
The model features green hydrogen technologies, covering its generation, storage, and reconversion to electricity. We assume that hydrogen cavern storage can be expanded without restrictions in every country. This is a deliberate simplification, as we aim to illustrate the relation between VRE portfolio droughts and long-duration electricity storage, but not to derive policy-relevant geographical allocations of hydrogen storage across Europe [caglayan_technical_2020, talukdar_techno-economic_2024]. We focus on hydrogen used for long-duration electricity storage and abstract from additional hydrogen demand of other sectors, hydrogen imports from other world regions, and hydrogen exchange between countries.
We investigate three scenarios with varying degrees of interconnection between the 33 countries. First, we model an island scenario without any cross-country exchange of electricity, i.e., domestic demand needs to be satisfied by domestic supply only. Second, we allow for unlimited exchange of electricity across countries, which represents the European copperplate scenario. Third, we complement these extreme scenarios by allowing for a more policy-relevant cross-border exchange of electricity according to the Ten-Year-Network-Development-Plan 2022 (scenario “Distributed Energy”) to prove the validity of the drought mass metric as a proxy for very severe VRE droughts.
Acknowledgments
We thank the entire research group “Transformation of the Energy Economy” at the German Institute for Economic Research (DIW Berlin) for valuable inputs and discussions, as well as conference participants of the EGU General Assembly 2022 and 2025, the International Energy Workshop 2022, the Conference on Climate, Weather and Carbon Risk in Energy and Finance 2022, the Next Generation Energy and Climate Workshop 2022, the International Conference Energy & Meteorology 2023, and the ENERDAY conference 2024 for valuable comments on earlier drafts. We acknowledge research grants by the Einstein Foundation (grant no. A-2020-612) and by the German Federal Ministry of Education and Research via the “Ariadne” projects (Fkz 03SFK5NO & 03SFK5NO-2).
Author Contributions
Martin Kittel: Conceptualization (lead), methodology, software, investigation (equal), data curation, visualization, writing - original draft, review and editing (equal). Wolf-Peter Schill: Conceptualization (support), investigation (equal), writing - review and editing (equal), project administration, funding acquisition.
[…]
SI Supplementary information
SI.1 Deployed thresholds
Figure SI.1 illustrates how the relative thresholds used in the search algorithm relate to absolute availability factors for each VRE technology and portfolio. Absolute thresholds are generally lower for solar PV than for wind power because of the lower average availability of solar PV.
(a) Solar PV.
(b) Onshore wind.
(c) Offshore wind.
(d) VRE portfolio.
Figure SI.1: Relative thresholds used for drought identification in absolute terms. The VRE portfolio thresholds are based on capacity-weighted composite time series.
SI.2 Additional illustrations of identified drought patterns
Figure SI.2 illustrates identified droughts lasting longer than one week or more, filtering out briefer events. This highlights very long-lasting below-average wind periods that may encompass multiple drought events of varying severity and solar seasonality. Combined in a portfolio, these periods are less severe (portfolio effect) and are further mitigated when assuming unconstrained geographical balancing (balancing effect). Figures SI.3SI.4, and SI.5 show identified drought patterns lasting at least one day for the years 2011/12, 1982/83, and 2013/14.
Figure SI.2: Identified drought patterns in 1996 and 1997 across all employed thresholds. For each technology-specific panel, a horizontal band indicates drought occurrences for the color-coded threshold of one country. To illustrate very persistent shortage situations, only droughts lasting longer than one week are displayed.
Figure SI.3: Identified drought patterns in 2011 and 2012 across all employed thresholds. For each technology-specific panel, a horizontal band indicates drought occurrences for the color-coded threshold of one country. To illustrate persistent patterns, only droughts lasting longer than one day are displayed. In Spain, low PV seasonality is accompanied by wind droughts, translating into portfolio droughts arising across the turn of the year 2011/12. In contrast, due to the exceptionally high wind availability in North Europe, no portfolio droughts occur across the turn of the year in these countries.
Figure SI.4: Identified drought patterns in 1982 and 1983 across all employed thresholds. For each technology-specific panel, a horizontal band indicates drought occurrences for the color-coded threshold of one country. To illustrate persistent patterns, only droughts lasting longer than one day are displayed. Except for Spain, no portfolio droughts occur at the beginning of a calendar year due to the absence of wind droughts in Northern Europe.
Figure SI.5: Identified drought patterns in 2013 and 2014 across all employed thresholds. For each technology-specific panel, a horizontal band indicates drought occurrences for the color-coded threshold of one country. To illustrate persistent patterns, only droughts lasting longer than one day are displayed. Brief wind droughts are infrequent in winter but longer-lasting, more severe, and more frequent in summer. Therefore, less severe portfolio droughts occur in the winter of 2013/14.
SI.3 Additional illustrations of frequency-duration distributions
Figure SI.6 illustrates the cumulative frequency-duration distributions for droughts lasting up to a full year for Germany, Spain, and the European copperplate scenario. Only higher thresholds identify droughts lasting longer than a few weeks. For thresholds $\tau<1$, portfolio droughts are generally briefer than single-technology droughts. To raise complementary insights on seasonality, Figure SI.7 shows seasonally differentiated distributions for the copperplate scenario. In general, longer droughts are more frequent in winter than in summer. Brief PV droughts are more frequent in summer and are detected by higher thresholds.
Figure SI.6: Example of cumulative frequency-duration distributions of drought events across all investigated thresholds $\tau\in[0.1,1]$ that may last up to one full year, sorting the frequencies of all events that are at least as long as a given duration. White space indicates the absence of droughts for given thresholds in the data. The contour lines represent the threshold-specific frequency.
Figure SI.7: Example of cumulative frequency-duration distributions of drought events across all investigated thresholds $\tau\in[0.1,1]$ under the assumption of unconstrained geographical balancing, sorting the frequencies of all events that are at least as long as a given duration. White space indicates the absence of droughts for given thresholds in the data. The contour lines represent the threshold-specific frequency. For illustration, the distributions are truncated at 360 hours, i.e., they show events with a maximum duration of just above two weeks. Frequencies of events lasting at least two days, one week, and a fortnight are marked for a relative threshold $\tau=0.75$.
SI.4 Return periods: most extreme events occur rarely
Return periods of VRE droughts are given by the reciprocal of yearly frequencies. The return period-duration distributions represent the right-hand side of the event distribution and indicate the longest event duration that can be expected to reoccur after a given number of years (Figure SI.8). Note that the return period monotonically increases and the maximum return period is limited by the 38 years of data that we investigate.
For each threshold, the maximum event duration increases with higher return periods (compare increasing contour lines and marked data points in Figure SI.8). Additionally, the return period-duration distributions are highly sensitive to the underlying threshold. For any given return period, higher thresholds lead to events with longer maximum durations, plateauing at 365 days for very high thresholds. Note that such high durations indicate below-average wind and solar years and should not be mistaken for actual drought events. The return period-duration distributions vary significantly between technologies and countries. For example, the maximum offshore wind drought for a relative threshold of $\tau=0.75$ in Germany that returns every 20 years lasts 198 days and is much shorter than the corresponding PV (289 days) and onshore wind droughts (297 days). By comparison, in Spain, such 20-year return period droughts are shorter for PV (216 days) and onshore wind (245 days) but longer for offshore wind (211 days).
When combined in a technology portfolio, the maximum duration decreases substantially for each return period in both Germany and Spain and also in the European copperplate, which is another instance of the portfolio effect described in the main body of this paper. The balancing effect is also visible: with perfect interconnection, the maximum drought duration further decreases due to a limited temporal correlation of wind droughts across Europe. In the case of solar PV, higher solar availability in South Europe during winter balances lower availability in Northern Europe.
Figure SI.8: Example of return period-duration distributions of rare drought events with an average annual frequency below 1 across all investigated thresholds $\tau\in[0.1,1]$. The contour lines along the return period axis represent the threshold-specific return period. The maximum duration of an event returning on average every 10, 20, and 30 years are marked for a threshold of $\tau=0.75$.
SI.5 Additional illustrations of maximum drought durations
Figure SI.9 shows the longest droughts obtained from the data for each year and threshold for Germany, Spain, and the European copperplate scenario. Figures SI.10 and SI.11 show the maximum duration of single drought events across all years for each country and threshold. Figure SI.12 illustrates the inter-annual variability of the maximum drought duration for each threshold and technology (portfolio).
Figure SI.9: Examples of most extreme duration of single drought events for each year and across all investigated thresholds $\tau\in[0.1,1]$. The contour lines represent threshold-specific maximum duration. White space indicates the absence of droughts for given thresholds in the data. The events with the highest and lowest duration across all years are marked for a threshold $\tau=0.75$. Arrows indicate values that are hidden in valleys of the distribution plane.
(a) Longest-lasting solar PV droughts. Maximum droughts duration in South European countries is lower than in Northern Europe. In the European copperplate scenario (CP), extreme durations in Northern Europe can be mitigated through geographical balancing with Southern Europe (balancing effect).
(b) Longest-lasting VRE portfolio droughts. For thresholds $\tau<1$, portfolio droughts are generally shorter than single-technology ones (portfolio effect).
Figure SI.10: Most extreme duration of single drought event of all years for each country and across all investigated thresholds $\tau\in[0.1,1]$. The latter are sorted in descending order from left to right according to the longest duration for a threshold $\tau=0.75$.
(a) Longest-lasting onshore wind droughts.
(b) Longest-lasting offshore wind droughts.
Figure SI.11: Most extreme duration of single drought event of all years for each country and across all investigated thresholds $\tau\in[0.1,1]$. The latter are sorted in descending order from left to right according to the longest duration for a threshold $\tau=0.75$. In the European copperplate scenario (CP), unconstrained geographical balancing mitigates the most extreme drought duration (balancing effect).
Figure SI.12: Most extreme duration of single drought events per year and across all investigated thresholds $\tau\in[0.1,1]$. Higher thresholds find long-lasting events. The year with the most extreme event duration varies across thresholds. The difference between years increases for increasing thresholds before it decreases again for very high thresholds. The ranking of years changes across thresholds. In general, combining technologies (portfolio effect) and countries (balancing effect) reduces the most extreme event duration.
SI.6 Additional illustrations of most extreme drought events
Figure SI.13 illustrates the most extreme drought events and optimal storage use in 1996/97 for selected regions assuming flat demand profiles, i.e., eliminating any diurnal or seasonal demand variability.
Figure SI.13: Identified most extreme drought events in 1996/97 occurring in winter (teal boxes). For countries in which the most extreme drought events occur in summer, these are additionally shown (gray boxes). For each region, portfolio drought occurrences lasting longer than one day for color-coded thresholds (upper panel) as well as exogenous flat demand profiles and optimized storage levels across three modeled interconnection scenarios (lower panel) are displayed. Note that the major storage discharging periods now perfectly coincide with the most extreme events identified by the drought mass metric in all countries, including the summer-time droughts in Sweden, Poland, and the United Kingdom.